# Boosting Prediction with Data Missing Not at Random

### v2

@@ 6.2 Evaluation Metrics
The first metric represents the sample-based maximum absolute error (S-MAE) (Chen and Yi,, 2024), defined as the infinity norm of the difference between the estimate and the true function with respect to the sample:
@@ 6.3 Analysis Methods
The first two methods, called R (reference) and N (naive), address two extreme scenarios: applying the conventional boosting procedure introduced in Section 2.2 to the full dataset $\mathcal{O}_{\text{full}}^{\text{TR}}$ and to the dataset $\mathcal{O}_{\text{missing}}^{\text{TR}}$, respectively. The next three methods apply boosting with modified loss functions described in Algorithm 1. Specifically, we employ the inverse propensity weight adjusted loss functions (4) with parametrically modeled $f(y|X=x,R=1)$ and nonparametrically estimated $f(y|X=x,R=1)$, respectively, denoted IPW and IPWN, where nonparametric estimation uses a Gaussian kernel with a rule-of-thumb bandwidth $h_{x}=(\sum_{i=1}^{n_{1}}r_{i})^{-\frac{1}{5}}\hat{\sigma}_{x}$, with $\hat{\sigma}_{x}$ denoting the sample standard deviation of the covariate, as suggested by Morikawa and Kim, (2021). Additionally, we consider the Buckley-James-type adjusted loss function (5), referred to as the BJ method, employing (15) with $N_{y}=20$.
@@
where $\eta$ is a transition point, which can be chosen as the $\alpha$th quantile of the set $\{|Y_{i}-f(X_{i})|:i=1,\ldots,n\}$, with $0\leq\alpha\leq 100$ (Friedman,, 2001). For the Huber loss, $\eta$ is iteratively specified as the 50th-quantile of $\{R_{i}|Y_{i}-f^{(m-1)}(X_{i})|:i=1,\ldots,n\}$ at iteration $m$, as in Friedman, (2001) and Bühlmann and Hothorn, (2007). As discussed by Bühlmann and Hothorn, (2007) and Hastie et al., (2009), the solution $f^{*}$ in (2) corresponds to median$(Y_{i}|X_{i})$ and mean$(Y_{i}|X_{i})$ for the $L_{1}$ loss and $L_{2}$ loss functions, respectively. Here “median$(Y_{i}|X_{i})$” and “mean$(Y_{i}|X_{i})$” represent the median and mean of the conditional probability function, $f_{j}(y|X_{i})$, of $Y_{i}$ given $X_{i}$, which is given by
@@
for Setting $j=1$ or 2, where $f_{j}(y|X_{i},R_{i}=0)$ and $f_{j}(y|X_{i},R_{i}=1)$ is defined in the Supplementary Materials. As shown in the Supplementary Materials, corresponding to Setting $j$ with $j=1,2$, mean$(Y_{i}|X_{i})$ equals
@@
In the MAR scenario, since $f_{j}(y|X_{i})=f_{j}(y|X_{i},R_{i}=1)=f_{j}(y|X_{i},R_{i}=0)$, which is identical to the probability density function (pdf) of (S.23) in the Supplementary Materials, median$(Y_{i}|X_{i})$ coincides with mean$(Y_{i}|X_{i})$ in (25). Given that the Huber loss behaves as the $L_{2}$ loss when $\left|Y_{i}-f(X_{i})\right|\leq\eta$, and as a linear form of the $L_{1}$ loss otherwise, the solution $f^{*}$ in (2) also corresponds to (25) if the Huber loss is used. However, in the MNAR scenario, median$(Y_{i}|X_{i})$ does not necessarily equal mean$(Y_{i}|X_{i})$, so $f^{*}$ generally lacks an analytic form when using the Huber loss.
@@ 6.4 Simulation Results
To visualize predicted values derived from different methods, Figure 3 shows boxplots of the average predicted values, ${n_{2}}^{-1}\sum^{n_{2}}_{i=1}\hat{f}^{*}_{n_{1}}(X_{i})$, across 500 simulations, along with the expected value $E(Y_{i})$, whose expression is provided in the Supplementary Materials. Under MAR, all methods yield similar results, though the N method tends to deviate more from the R method than the proposed methods. However, with MNAR in Setting 1, the R and proposed methods (i.e., IPW, IPWN, and BJ) produce fairly close results, while the N methods exhibit noticeable bias. In Setting 2, with a large number of correlated covariates, IPWN does not perform satisfactorily, as the MNAR estimation procedure of Morikawa and Kim, (2021) assumes independent covariates.
@@
Finally, we assess the convergence of the proposed boosting methods. The results confirm the convergence of our boosting algorithm, as long as estimates of $\pi_{2}(y,x)$, defined in the Supplementary Materials, using the method of Morikawa and Kim, (2021) are available; further details are provided in the Supplementary Materials. We note that the method proposed by Morikawa and Kim, (2021) is specifically designed for MNAR data and may not perform well for MAR data, particularly in scenarios with high-dimensional covariates. In such cases, divergence can occur frequently, as acknowledged by Morikawa and Kim, (2021), thus preventing the estimation of $\pi_{2}(y,x)$ and halting the boosting procedure. This issue may stem from various factors related to starting values for the nonlinear optimization function, the complexity of the optimization process, and the large number of parameters.
@@ 7 Analysis of KLIPS Data
To illustrate the proposed methods, we apply them to analyze a dataset from the Korean Labor and Income Panel Survey (KLIPS), which contains demographics information for 2501 Korean workers. The KLIPS data are frequently analyzed in the MNAR literature, including Kim and Yu, (2011), Wang et al., (2014), Shao and Wang, (2016), Morikawa et al., (2017), Morikawa and Kim, (2021), and Bian et al., 2025b .
@@
Assume that the missing data indicator $R_{i}$ follows the Bernoulli distribution, Bernoulli($\pi(Y_{i},X_{i})$), with
@@
where $\gamma_{j}$ is the parameter for $j=0,1,\ldots,4$. Given the missing indicator $R_{i}=1$ and covariates $X_{i,1}$, $X_{i,2}=j$, $X_{i3}=k$ for $j=1,2,3$ and $k=1,2$, the response $Y_{i}$ follows $N(\beta_{0,j,k}+\beta_{1,j,k}X_{i,1}+\beta_{2,j,k}X_{i,1}^{2}+\beta_{3,j,k}X_{i,1}^{3}+\beta_{4,j,k}X_{i,1}^{4},\sigma_{j,k}^{2})$, with parameters $\beta_{j,k}=(\beta_{0,j,k},\beta_{1,j,k},\beta_{2,j,k},\beta_{3,j,k},\beta_{4,j,k})^{\top}$ and $\sigma_{j,k}$. Following Morikawa and Kim, (2021), for each combination of $j$ and $k$, we use the stepwise AIC to select the best model.

### v3

@@ 6.2 Evaluation Metrics
The first metric represents the sample-based maximum absolute error (S-MAE) (Chen and Yi, 2024), defined as the infinity norm of the difference between the estimate and the true function with respect to the sample:
@@ 6.3 Analysis Methods
The first two methods, called R (reference) and N (naive), address two extreme scenarios: applying the conventional boosting procedure introduced in Section 2.2 to the full dataset $\mathcal{O}_{\text{full}}^{\text{TR}}$ and to the dataset $\mathcal{O}_{\text{missing}}^{\text{TR}}$, respectively. The next three methods apply boosting with modified loss functions described in Algorithm 1. Specifically, we employ the inverse propensity weight adjusted loss functions (4) with parametrically modeled $f(y|X=x,R=1)$ and nonparametrically estimated $f(y|X=x,R=1)$, respectively, denoted IPW and IPWN, where nonparametric estimation uses a Gaussian kernel with a rule-of-thumb bandwidth $h_{x}=(\sum_{i=1}^{n_{1}}r_{i})^{-\frac{1}{5}}\hat{\sigma}_{x}$, with $\hat{\sigma}_{x}$ denoting the sample standard deviation of the covariate, as suggested by Morikawa and Kim (2021). Additionally, we consider the Buckley-James-type adjusted loss function (5), referred to as the BJ method, employing (15) with $N_{y}=20$.
@@
[…]
@@
[…]
@@
[…]
@@ 6.4 Simulation Results
To visualize predicted values derived from different methods, Figure 3 shows boxplots of the average predicted values, ${n_{2}}^{-1}\sum^{n_{2}}_{i=1}\hat{f}^{*}_{n_{1}}(X_{i})$, across 500 simulations, along with the expected value $E(Y_{i})$, whose expression is provided in the Supplementary Materials. Under MAR, all methods yield similar results, though the N method tends to deviate more from the R method than the proposed methods. However, with MNAR in Setting 1, the R and proposed methods (i.e., IPW, IPWN, and BJ) produce fairly close results, while the N methods exhibit noticeable bias. In Setting 2, with a large number of correlated covariates, IPWN does not perform satisfactorily, as the MNAR estimation procedure of Morikawa and Kim (2021) assumes independent covariates.
@@
Finally, we assess the convergence of the proposed boosting methods. The results confirm the convergence of our boosting algorithm, as long as estimates of $\pi_{2}(y,x)$, defined in the Supplementary Materials, using the method of Morikawa and Kim (2021) are available; further details are provided in the Supplementary Materials. We note that the method proposed by Morikawa and Kim (2021) is specifically designed for MNAR data and may not perform well for MAR data, particularly in scenarios with high-dimensional covariates. In such cases, divergence can occur frequently, as acknowledged by Morikawa and Kim (2021), thus preventing the estimation of $\pi_{2}(y,x)$ and halting the boosting procedure. This issue may stem from various factors related to starting values for the nonlinear optimization function, the complexity of the optimization process, and the large number of parameters.
@@ 7 Analysis of KLIPS Data
To illustrate the proposed methods, we apply them to analyze a dataset from the Korean Labor and Income Panel Survey (KLIPS), which contains demographics information for 2501 Korean workers. The KLIPS data are frequently analyzed in the MNAR literature, including Kim and Yu (2011), Wang et al. (2014), Shao and Wang (2016), Morikawa et al. (2017), Morikawa and Kim (2021), and Bian et al. (2025b).
@@
[…]
@@
where $\gamma_{j}$ is the parameter for $j=0,1,\ldots,4$. Given the missing indicator $R_{i}=1$ and covariates $X_{i,1}$, $X_{i,2}=j$, $X_{i3}=k$ for $j=1,2,3$ and $k=1,2$, the response $Y_{i}$ follows $N(\beta_{0,j,k}+\beta_{1,j,k}X_{i,1}+\beta_{2,j,k}X_{i,1}^{2}+\beta_{3,j,k}X_{i,1}^{3}+\beta_{4,j,k}X_{i,1}^{4},\sigma_{j,k}^{2})$, with parameters $\beta_{j,k}=(\beta_{0,j,k},\beta_{1,j,k},\beta_{2,j,k},\beta_{3,j,k},\beta_{4,j,k})^{\top}$ and $\sigma_{j,k}$. Following Morikawa and Kim (2021), for each combination of $j$ and $k$, we use the stepwise AIC to select the best model.
@@ 8 Discussion
Supplementary Materials
Text document:
A .pdf file contains regularity conditions, proofs of Propositions 1 and 2, Theorems 1 and 2, and additional simulation results.
R-code:
A .zip file, named “code.zip”, includes R scripts for the simulations and data analyses presented in Sections 6 and 7.
Acknowledgments
The authors thank the review team for their comments on the initial version of the manuscript and Dr. Morikawa for providing the KLIPS data and the R code used in Morikawa and Kim (2021). Yi is a Canada Research Chair in Data Science (Tier 1). Her research is supported by the Natural Sciences and Engineering Research Council of Canada (NSERC) and the Canada Research Chairs Program.